# Vision-Based Power Line Cables and Pylons Detection for Low Flying Aircrafts

### v1

@@ 4.4 Performance Evaluation
Table 1: Comparison of power line cables detection performance between our method and DeepWireCNN (DWCNN) [1] on the NE-VBW and DDLN datasets. Results for the two size variants of our models are shown – DDLN-S and DDLN-L. DDLN-S has 2.7 million trainable parameters, about the same amount as in the DeepWireCNN model.
@@
Table 1 shows the main performance comparison between the two variants of our model and for the DeepWireCNN [1], both on the NE-VBW dataset and our DDLN dataset. We also measured inference speed as a number of processed $1\times 3\times 4096\times 3000$ random input tensors per second (i.e. frames per second, FPS) on a NVidia V100 GPU. Additionally, Figure 4 depicts qualitative results. After tuning the hyperparameters of the DeepWireCNN model we achieved 0.70$\pm$0.09 AP for the segmentation of cables at coarse resolution and the values for the remaining metrics as reported in the first row of the Table 1. The results on NE-VBW dataset demonstrate that both the small and the large variant of our model outperform DeepWireCNN on all cable detection metrics. Additionally, they do so without increasing the inference time. We also observed that our models achieve high performance scores earlier in the learning process than the DeepWireCNN. We attribute this to the initialization of the networks with the pretrained weights. Focusing only on our models, it is also clear that the larger variant achieves better or equivalent performance to its smaller counterpart across all cable detection metrics.
Figure 5: Exemplary predictions of the two size variants of our model trained jointly for cable and pole detection on the DDLN dataset. The inputs are $1024\times 1024$ px crops from full-resolution images from the validation set. Predictions were generated by models initialized from checkpoints at epochs with highest validation cable segmentation quality metric.
@@
In Table 2 we compare the cable detection performance of our DDLN-M model against that of PIDNet-S [2]. As the latter is designed for real-time performance, we also report the model size and inference speeds for all networks. The size is given as the number of millions of trainable parameters in the respective model. Regarding detection performance, our DDLN-M model, which has the same size as PIDNet-S, yields slightly better cable detection quality and a similar F1 score. Given this result, we conducted additional experiments with PIDNet-S trained to simultaneously detect cables and pylons, as described in section 4.2, and compared its performance on both tasks with our joint detection model, which we detail below.
Table 3: Comparison of joint cables and poles detection performance on DDLN dataset between DDLN-M and PIDNet-S [2]. PIDNet-S was trained for cable and pole detection jointly as described in section 4.2. The reported metrics are measured at the epochs with the highest validation cable detection quality.
@@
To continue the comparison of our method with the PIDNet-S [2] baseline, we added to the latter a second segmentation output head to enable it to also perform pylon detection. We report our results in Table 3 and shown in Figure 6 the evolution of the validation score during training. Our DDLN-M model still outperforms PIDNet-S on the cable detection quality, as on the previously equal F1 score. It also delivers minimally higher pylon detection quality, while it records a slightly lower F1 score than PIDNet-S. Identical observations can be made by looking at the Figure 6 where validation performance curves of our DDLN-M model are consistently, and across all validated epochs, above those of PIDNet-S for cable detection quality and F1 score. As in the metrics reported in Table 3, pylon detection quality of DDLN-M is also minimally above that of PIDNet-S, while pylon detection F1 score is slightly below the one of the segmentation baseline. Furthermore, recall from Table 1 that DDLN-L performs even better.
Regarding the network inference speeds, PIDNet-S is the faster one with 22.2 FPS compared to our 10.7. However, this high speed provided by PIDNet-S may not be absolutely necessary for a possible deployment of our method. We have successfully integrated DDLN-S, the less performant but faster smaller alternative of our model, in a hardware system operating in real-time onboard of a helicopter. We run the system in flight and reached 5 FPS inference speed. This is already a good running speed for the application purpose of an aircraft safety system.
@@
Figure 6: Evolution of validation metrics – quality and F1 for cable and pylon detection – in 5-fold cross validation runs reported in table 3 for DDLN-M and PIDNet-S [2], both trained for joint detection task. The solid lines are means of respective metrics, while the shaded area correspond to standard deviation.
@@
Finally, we compare our models against RT-DETR [3] in table 4 focusing on the pole detection alone. In the presented results we adapt two kinds of thresholds used to turn RT-DETR predictions into distance masks: a cutoff confidence threshold filtering out the least certain bounding box predictions and a distance mask threshold used to convert distance mask, derived from predicted bounding boxes, into a segmentation mask. The choice of these hyperparameters is different depending on whether we optimize for best quality or F1 score, therefore we report both underlining the metric with respect to which we optimized them.
@@
One reason for our superior performance on the CCQ metrics is due to the fact that those give, by design, more favourable evaluation to the method inherently using distance mask regression than to the one in which the distance masks are only derived from a different representation – precisely regressed bounding boxes, as done for RT-DETR. The distance mask regression objective imposes a distance-dependent structure on the outputs regressed by our networks. This structure is also maintained when a segmentation mask is obtained through distance mask thresholding. Finally, the CCQ metrics, which relax standard precision and recall, also have a distance-dependent component in their use of error tolerance margin, which interplays favourably with the imposed distance-dependent structure of our targets. Meanwhile, the distance masks generated for RT-DETR are only derived from originally predicted bounding boxes and therefore do not inherit that structure in the resulting distance mask representation of the object detector’s output.
@@
Table 4: Comparison of pole detection on DDLN dataset between our distance mask regression method and RT-DETR [3] combined with SAHI [74]. Our models were trained jointly for cables and poles detection. The metrics for our models are reported for epochs with the best mean validation pylon detection quality. For RT-DETR runs, the reported epochs were selected as ones with maximal mean validation value of the underlined metric – quality or F1. The outputs of the models are evaluated with distance masks of 32 times smaller resolution than the input image size, corresponding to the size of the output of the DDLN-H model.
@@ 4.5 Implementation Details
Distance and segmentation masks. To generate learning targets for our models, we converted the annotations into distance masks by computing unsigned Euclidean distance transform to the detected objects (cables and pylons). For the DDLN, we treat as object areas 5-pixel-thick regions around annotation polylines and the insides of the pylons’ bounding boxes. The target distance regression masks are first computed in pixel distance units, then clamped to the maximum distance value of 128 and finally normalized to the $[0,1]$ range. The distance masks are downsampled through $16\times 16$ min-pooling operation to provide the targets for learning at the coarse resolution. To generate segmentation masks used by our baselines – DeepWireCNN [1] and PIDNet-S [2], we binarized full resolution distance masks with a distance value threshold of 32. Then we downsampled the labels to match the output resolution. To perform the downsampling $n$ times we look at the center 4 pixels of $n\times n$ patches and if they contain a foreground object we label the pixel in the downsampled output as foreground as well, and as background otherwise.
Patch sampling. For the experiments run on NE-VBW dataset and for DeepWireCNN, we followed the patch sampling procedure proposed by Stambler et al. [1]. For all other experiments, we use our own sampling method. We select patch center locations dynamically during training from the high-resolution images, based on the data annotations. Patches are sampled as centered on some part of the target detected object or within its vicinity, with a set maximal allowed distance to the closest object and ensuring that the entire patch remains within the image boundaries. This effectively implements sampling position with random horizontal and vertical jitter. When training our cable detection and joint cable and pylon detection models we do not sample purely background patches. We did so only when training the baselines models, where this setting was one of the hyperparameters. In most experiments we centered the sampled patches on power line cables. This procedure was different when training pylon only detection model – RT-DETR – where we analogously centered the sampled patches on pylons instead of cables. It was also different when training PIDNet-S [2] and DDLN-M models for comparing the joint detection performance in table 3. In those experiments we centered patches within the union of close-to-cables and close-to-pylons areas.
@@
Hyperparameters. For all models we compare against, we tuned selected sets of their hyperparameters on one of the splits of the NE-VBW dataset or on our train-validation development split of DDLN. For this procedure we used SMAC3 [88], a framework implementing Bayesian optimization and Hyperband (BOHB) [89] for hyperparameter optimization. To obtain the reported performance metrics values we run the 5-fold cross validation with the model and data augmentations configured according to the found best performing hyperparameters configurations. Finally, when thresholding on the probabilities predicted by DeepWireCNN [1] and PIDNet-S [2] models, as well as on confidence scores and on distance masks derived from RT-DETR [3] predictions, we also choose threshold values which maximize the validation metrics of the obtained target object segmentation. We choose these threshold values for each data fold separately, maximizing their respective validation performance, therefore the reported mean metrics represent the best achievable performance, providing very strong baseline results for our models to compare with.
For our cable detection models, we selected the hyper parameters firstly mostly manually, throughout the method development process, and then by trying out several configurations and running with them the complete 5-fold cross validation experiments. Finally, we selected the configurations achieving the best maximal mean validation cable segmentation quality. We also selected thresholds, same for all folds in a single 5-fold cross validation run, for converting predicted distance masks into segmentations so as to maximize the detection validation quality metric, both for cables and pylons. We resorted to some automated tuning with SMAC3 [88] for our comparison with PIDNet-S [2] for joint cables and pylons detection.
@@ 4.6 Ablations
Table 5: Ablation study demonstrating the impact on detection performance of the components in our proposed method. The ablation experiments have been conducted with the DDLN-S model on the NE-VBW [1] dataset. “Pretr.” indicates the initializing the model parameters with the weights of ConvNext-Tiny pretrained on the ImageNet [43] dataset. “LIF” refers to logarithmically scaled inverse frequency weights in $\mathcal{L}_{\rm dat}$. $\mathcal{L}_{\texttt{MALIS}}$ [51] referes to the loss term promoting connectivity in cable detections.
@@
For our complete setup we used both the $\mathcal{L}_{\rm dat}$ and $\mathcal{L}_{\texttt{MALIS}}$, which is the configuration for the model presented in all result tables apart from table 5. However, comparing the evolution of the validation cable segmentation quality metric during training for the complete setup and for those using either of the loss function components in isolation, it is apparent that these configurations are actually leading to equal performance. We also note that the final scores depend here on the value of the used binarization threshold with which we convert distance masks into segmentation masks. In our experiments we observed that with a lower threshold value, using solely $\mathcal{L}_{\texttt{MALIS}}$ component gives better cable segmentation quality than $\mathcal{L}_{\rm dat}$ alone or both loss components jointly. However, overall the performance was lower than in the ablation experiments presented here in which we used a larger binarization threshold. Ultimately, for all the runs of our models reported in the previously described result tables we used both loss components and the larger binarization threshold value.
@@ 5 Conclusion
We have proposed a vision-based approach to joint detection of power line cables and transmission towers as seen by low-flying aircrafts. To this end, we have combined a modern convolutional architecture with transfer learning and a loss function adapted to curvilinear structure delineation Oner et al. [51]. We use a single network for both detection tasks and demonstrated its performance on two benchmarking datasets.

### v3

@@ 4.4 Performance Evaluation
Table 1: Comparison of power line cables detection performance between our method and DeepWireCNN (DWCNN) [1] on the NE-VBW and DDLN datasets. Results for the two size variants of our models are shown – DDLN-S and DDLN-L. DDLN-S has 2.7 million trainable parameters, about the same amount as the DeepWireCNN model.
@@
Table 1 shows the main performance comparison between the two variants of our model and the DeepWireCNN [1], both on the NE-VBW dataset and our DDLN dataset. We also measured inference speed as a number of processed $1\times 3\times 4096\times 3000$ random input tensors per second (i.e. frames per second, FPS) on a NVidia V100 GPU, using automatic mixed-precision arithmetic. Additionally, Figure 4 depicts qualitative results. After tuning the hyperparameters of the DeepWireCNN model we achieved 0.70$\pm$0.09 AP for the segmentation of cables at coarse resolution and the values for the remaining metrics as reported in the first row of the Table 1. The results on NE-VBW dataset demonstrate that both the small and the large variant of our model outperform DeepWireCNN on all cable detection metrics. Additionally, they do so without increasing the inference time. We also observed that our models achieve high performance scores earlier in the learning process than the DeepWireCNN. We attribute this to the initialization of the networks with the pretrained weights. Focusing only on our models, it is also clear that the larger variant achieves better or equivalent performance to its smaller counterpart across all cable detection metrics.
Figure 5: Exemplary predictions of the two size variants of our model trained jointly for cable and pylon detection on the DDLN dataset. The inputs are $1024\times 1024$ px crops from full-resolution images from the validation set. Predictions were generated by models initialized from checkpoints at epochs with highest validation cable segmentation quality metric.
@@
In Table 2 we compare the cable detection performance of our DDLN-M model against that of PIDNet-S [84]. As the latter is designed for real-time performance, we also report the model size and inference speeds for the networks. The size is given as the number of millions of trainable parameters in the respective model. Regarding detection performance, our DDLN-M model, which has the same size as PIDNet-S, yields slightly better cable detection quality and a similar F1 score. Given this result, we conducted additional experiments with PIDNet-S trained to simultaneously detect cables and pylons, as previously described in section 4.2, and compared its performance on both tasks with our joint detection model, which we detail below.
Table 3: Comparison of joint cables and pylons detection performance on DDLN dataset between DDLN-M and PIDNet-S [84]. PIDNet-S was trained for cable and pylon detection jointly as described in section 4.2. The reported metrics are measured at the epochs with the highest validation cable detection quality.
@@
To continue the comparison of our method with the PIDNet-S [84] baseline, we added to the latter a second segmentation output head to enable it to also perform pylon detection. We report our results in Table 3 and shown in Figure 6 the evolution of the validation score during training. Our DDLN-M model still outperforms PIDNet-S on the cable detection quality, as well as on the previously equal F1 score. It also delivers minimally higher pylon detection quality, while it records a slightly lower F1 score than PIDNet-S. Identical observations can be made by looking at the Figure 6 where validation performance curves of our DDLN-M model are consistently, and across all validated epochs, above those of PIDNet-S for cable detection quality and F1 score. As in the metrics reported in Table 3, pylon detection quality of DDLN-M is also minimally above that of PIDNet-S, while pylon detection F1 score is slightly below the one of the segmentation baseline. Furthermore, recall from Table 1 that DDLN-L performs even better.
Regarding the network inference speeds, PIDNet-S is the faster one with 22.2 FPS compared to our 10.7. However, this high speed provided by PIDNet-S may not be absolutely necessary for a possible deployment of our method. We have successfully integrated DDLN-S, the less performant but faster smaller alternative of our model, in the hardware system operating in real-time onboard of a helicopter, as described in section 3.3. We run the system in flight and reached average inference speed of 5 FPS. This is already a good running speed for the application purpose of an aircraft safety system.
@@
Figure 6: Evolution of validation metrics – quality and F1 for cable and pylon detection – in 5-fold cross validation runs reported in table 3 for DDLN-M and PIDNet-S [84], both trained for joint detection task. The solid lines are means of respective metrics, while the shaded areas correspond to standard deviations.
@@
Finally, we compare our models against RT-DETR [27] in table 4 focusing on the pylon detection alone. In the presented results we adapt two kinds of thresholds used to turn RT-DETR predictions into distance masks: a cutoff confidence threshold filtering out the least certain bounding box predictions and a distance mask threshold used to convert a distance mask, derived from predicted bounding boxes, into a segmentation mask. The choice of these hyperparameters is different depending on whether we optimize for best quality or F1 score, therefore we report both, underlining the metric with respect to which we optimized the thresholds.
@@
One reason for our superior performance on the CCQ metrics is due to the fact that those give, by design, more favourable evaluation to the method inherently using distance mask regression than to the one in which the distance masks are only derived from a different representation – precisely regressed bounding boxes, as done for RT-DETR. The distance mask regression objective imposes a distance-dependent structure on the outputs regressed by our networks. This structure is also maintained when a segmentation mask is obtained through distance mask thresholding. Finally, the CCQ metrics, which relax standard precision and recall, also have a distance-dependent component in their use of error tolerance margin, which interplays favourably with the imposed distance-dependent structure of our targets. Meanwhile, the distance masks generated for RT-DETR are only derived with exact Euclidean distance transform from originally predicted bounding boxes. Therefore, they do not naturally inherit the distance-dependent structure in the resulting distance mask representation of the object detector’s original output.
@@
Table 4: Comparison of pylon detection on DDLN dataset between our distance mask regression method and RT-DETR [27] combined with SAHI [73]. Our models were trained jointly for cables and pylons detection. The metrics for our models are reported for epochs with the best mean validation pylon detection quality. For RT-DETR runs, the reported epochs were selected as ones with maximal mean validation value of the underlined metric – quality or F1. The outputs of the models are evaluated with distance masks of 32 times smaller resolution than the input image size, corresponding to the size of the output of the DDLN-H model.
@@ 4.5 Implementation Details
Distance and segmentation masks. To generate learning targets for our models, we converted the annotations into distance masks by computing unsigned Euclidean distance transform to the detected objects (cables and pylons). For the DDLN, we treat as object areas 5-pixel-thick regions around annotation polylines and the insides of the pylons’ bounding boxes. The target distance regression masks are first computed in pixel distance units, then clamped to the maximum distance value of 128 and finally normalized to the $[0,1]$ range. The distance masks are downsampled through $16\times 16$ min-pooling operation to provide the targets for learning at the coarse resolution.
To generate segmentation masks used by our baselines – DeepWireCNN [1] and PIDNet-S [84], we binarized full resolution distance masks with a distance value threshold of 32. Then we downsampled the labels to match the output resolution. To perform the $n$ times downsampling of the segmentation masks we look at the center 4 pixels of $n\times n$ patches and if they contain a foreground object we label the pixel in the downsampled output as foreground as well, and as background otherwise.
Patch sampling. For the experiments run on NE-VBW dataset and also for those using DeepWireCNN, we followed the patch sampling procedure proposed by Stambler et al. [1]. For all other experiments, we use our own sampling method. We select patch center locations dynamically during training from the high-resolution images, based on the data annotations. Patches are sampled as centered on some part of the target detected object or within its vicinity, with a set maximal allowed distance to the closest object and ensuring that the entire patch remains within the image boundaries. This effectively implements sampling position with random horizontal and vertical jitter. When training our cable detection and joint cable and pylon detection models we do not sample purely background patches. We did so only when training the baselines models, where this setting was one of the hyperparameters. In most experiments we centered the sampled patches on power line cables. This procedure was different when training pylon-only detection model – RT-DETR [27] – where we analogously centered the sampled patches on pylons instead of cables. It was also different when training PIDNet-S [84] and DDLN-M models for comparing the joint detection performance in table 3. In those experiments we centered patches within the union of close-to-cables and close-to-pylons areas.
@@
Hyperparameters. For all models we compare against, we tuned selected sets of their hyperparameters on one of the splits of the NE-VBW dataset or on our train/validation development split of DDLN. For this procedure we used SMAC3 [90], a framework implementing Bayesian optimization and Hyperband (BOHB) [91] for hyperparameter optimization. To obtain the reported performance metrics values we run the 5-fold cross validation with the model and data augmentations configured according to the found best performing hyperparameters configurations. Finally, when thresholding on the probabilities predicted by DeepWireCNN [1] and PIDNet-S [84] models, as well as on confidence scores and on distance masks derived from RT-DETR [27] predictions, we also choose threshold values which maximize the validation metrics of the obtained target object segmentation. We choose these threshold values for each data fold separately, maximizing their respective validation performance, therefore the reported mean metrics represent the best achievable performance, providing very strong baseline results for our models to compare with.
For our cable detection models, we selected the hyper parameters firstly mostly manually, throughout the method development process, and then by trying out several configurations and running with them the complete 5-fold cross validation experiments. Finally, we selected the configurations achieving the best maximal mean validation cable or pylon segmentation quality. We also selected thresholds, same for all folds in a single 5-fold cross validation run, for converting predicted distance masks into segmentations so as to maximize the detection validation quality metric, both for cables and pylons. We resorted to some automated tuning with SMAC3 [90] for our comparison with PIDNet-S [84] for joint cables and pylons detection.
@@ 4.6 Ablations
Table 5: Ablation study demonstrating the impact of the components in our proposed method on cable detection performance. The ablation experiments have been conducted with the DDLN-S model on the NE-VBW [1] dataset. “Pretr.” indicates the initializing the model parameters with the weights of ConvNeXt-Tiny pretrained on the ImageNet [42] dataset. “LIF” refers to logarithmically scaled inverse frequency weights in $\mathcal{L}_{\rm dat}$. $\mathcal{L}_{\texttt{MALIS}}$ [50] referes to the loss term promoting connectivity in cable detections.
@@
For our complete setup we used both the $\mathcal{L}_{\rm dat}$ and $\mathcal{L}_{\texttt{MALIS}}$, which is the configuration for the model presented in all result tables apart from table 5. However, comparing the evolution of the validation cable segmentation quality metric during training for the complete setup and for those using either of the loss function components in isolation, it is apparent that these configurations are actually leading to equal performance for cable-only detection. We also note that the final scores depend here on the value of the used binarization threshold with which we convert distance masks into segmentation masks. In our experiments we observed that with a lower threshold value, using solely $\mathcal{L}_{\texttt{MALIS}}$ component gives better cable segmentation quality than $\mathcal{L}_{\rm dat}$ alone or both loss components jointly. However, overall the performance was lower than in the ablation experiments presented here in which we used a larger binarization threshold. Ultimately, for all the runs of our models reported in the previously described result tables we used both loss components and the larger binarization threshold value.
@@ 5 Conclusion
We have proposed a vision-based approach to joint detection of power line cables and transmission towers as seen by low-flying aircraft. To this end, we have combined a modern convolutional architecture with transfer learning and a loss function adapted to curvilinear structure delineation [50]. We have used a single network for both detection tasks and have demonstrated its performance on two benchmarking datasets.
@@
Declarations
Funding
This work was supported under an Innosuisse Grant number 57140.1 IP-ICT funding the collaboration between Daedalean AG and EPFL.
Competing Interests
Employment and financial interests Ad Eberhard, Ivan Egorov, Philipp Krüsi and Grigory Yakushev declare that this work was done as part of their employment with Daedalean AG. Jakub Gwizdała, Doruk Oner, Soumava Kumar Roy, Mian Akbar Shah and Pascal Fua declare they have no financial interests.
Non-financial interests None.
Dual-use
The authors declare that the solutions and results of this work are intended for and were developed for commercial/industrial civilian aircraft operations. They shall not be used in any context in which any harm done to human lives, property and/or infrustructure as well as to natural habitats and ecosystems is a direct or an indirect consequence of their use.
Data
This work used publicly available Near Earth Autonomy Vision Based Wire Detection (NE-VBW) dataset which is described at https://sites.google.com/nearearth.aero/ne-vbwd and can be downloaded from https://drive.google.com/drive/folders/1A4hViJWaf9mwSISBAG629CGg_VnVHRtb. Used 5-fold cross validation split of this dataset content can be accessed at https://drive.google.com/file/d/1eodtX6bU-6lBSz1-rlIL5ptocu8kqV3t/view?usp=sharing. The collected DDLN dataset remains proprietary to Daedalean AG and is not publicly available.
Authors’ Contribution Statements
Conceptualization: Grigory Yakushev, Pascal Fua; Funding acquisition: Pascal Fua; Data preparation: Ad Eberhard, Grigory Yakushev; Methodology: Doruk Oner, Pascal Fua, Grigory Yakushev; Investigation: Jakub Gwizdała, Mian Akbar Shah, Soumava Kumar Roy; Project administration: Pascal Fua, Ad Eberhard, Grigory Yakushev; Software: Jakub Gwizdała, Ivan Egorov, Mian Akbar Shah, Doruk Oner; Supervision: Grigory Yakushev, Pascal Fua, Doruk Oner; Writing – original drafts: Jakub Gwizdała, Ivan Egorov, Pascal Fua; Writing – review & editing: Pascal Fua, Doruk Oner, Ivan Egorov, Philipp Krüsi, Soumava Kumar Roy.